OpenStudy (anonymous):

How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Where is the equation?

OpenStudy (anonymous):

4(2x+3)=2(2x-4)

OpenStudy (anonymous):

Distribute first

OpenStudy (anonymous):

ok ik the answer will come out as x=-5

OpenStudy (anonymous):

There is only one answer, but it's not x=-5

OpenStudy (anonymous):

4(2x+3) = 8x+12 2(2x+4) = 4x+8 8x+12=4x+8 4x+12=8 4x = -4 x = -1

OpenStudy (anonymous):

4(2x+3)=2(3x-4) <<<<< this is the right problem sorry i put the problem in wrong the first time